The difference in returns between Treasury bills and the FTSE Bursa Malaysia (FBM) KLCI is 5.50%. The Treasury bill rate is 3% . Krogh Enterprise has a beta of 1.60. What's the market return?

What is the required rate of return of Krogh Enterprise?


1
Expert's answer
2022-03-30T13:01:01-0400
market return is;

5.50 + 3=8.50%


What is the required rate of return of Krogh Enterprise?


(5.5\times 1.60) +3=11.80%